OpenStudy (anonymous):

A woman 5 ft tall walks at the rate of 7.5 ft/sec away from a streetlight that is 10 ft above the ground. At what rate is the tip of her shadow moving? At what rate is her shadow lengthening?

OpenStudy (caominhim):

angle with horizontal to top of pole = \[ \tan^{-1}(\frac{poleHeight}{current position})\] distance from pole to tip of shadow = \[\frac{poleHeight}{\tan(\angle horz)}\] current shadow length = tip of shadow - current position rate of shadow increase =\[\frac{d}{dx} currentShadowLength\] x being current position of the person rate that the tip is moving is\[\frac{d}{dx} distanceFrmoPoleToTipOfShadow\] x also being current position

OpenStudy (anonymous):

Your explanation is confusing to me..

OpenStudy (caominhim):

have you learned derivatives?

OpenStudy (anonymous):

yes

OpenStudy (caominhim):

so what part is confusing?

OpenStudy (anonymous):

how to find the solution

OpenStudy (caominhim):

follow the steps i just gave you, it's a series of equations you need to put together